# Set the Clock: Temporal Alignment of Pretrained Language Models

### v1

@@ Abstract
Language models (LMs) are trained on web text originating from many points in time and, in general, without any explicit temporal grounding. This work investigates the temporal chaos of pretrained LMs and explores various methods to align their internal knowledge to a target time, which we call “temporal alignment.” To do this, we first automatically construct a dataset containing 20K time-sensitive questions and their answers for each year from 2000 to 2023. Based on this dataset, we empirically show that pretrained LMs (e.g., LLaMa2), despite having a recent pretraining cutoff (e.g., 2022), mostly answer questions using earlier knowledge (e.g., in 2019). We then develop several methods, from prompting to finetuning, to align LMs to use their most recent knowledge when answering questions, and investigate various factors in this alignment. Our experiments show that aligning LLaMa2 to the year 2022 can boost its performance by up to 62% relatively as measured by that year, even without mentioning time information explicitly, indicating the possibility of aligning models’ internal sense of time after pretraining. Finally, we find that alignment to a historical time is also possible, with up to 2.8×2.8$\times$2.8 × the performance of the unaligned LM in 2010 if finetuning models to that year. These findings hint at the sophistication of LMs'  internal knowledge organization and the necessity of tuning them properly. [^1] Our dataset and code will be released at https://github.com/yizhongw/llm-temporal-alignment.
@@
Bowen Zhaothanks: Equal ${}^{\spadesuit}$     Zander Brumbaughfootnotemark: ${}^{\spadesuit}$     Yizhong Wangfootnotemark: ${}^{\spadesuit\clubsuit}$ Hannaneh ${}^{\spadesuit\clubsuit}$      Noah A. ${}^{\spadesuit\clubsuit}$$^{\spadesuit}$niversity of Washington   ${}^{\clubsuit}$llen Institute for AI bowen98@uw.edu   {brumbzan, yizhongw}@cs.washington.edu
@@ 1 Introduction
We first introduce our TemporalAlignmentQA dataset (TAQA, §2) consisting of 20,148 questions, each with at least five different answers between 2000 and 2023. This dataset is automatically constructed based on Wikipedia tables that contain temporal information, making it easily scalable for future study as the world keeps changing. Complementary to previous temporal QA datasets (Chen et al., 2021; Dhingra et al., 2022; Kasai et al., 2022; Liska et al., 2022; Tan et al., 2023; Wei et al., 2023), TAQA focuses on facts that changed multiple times during a recent period when most pretraining data is collected, so that it can be used to probe the temporal knowledge distribution of pretrained LMs. Figure 1 plots the F1 score of several representative LMs relative to ground-truth answers from each year, demonstrating that they tend to use earlier knowledge to answer questions, even when they have a very recent pretraining cutoff date (e.g., performance peaks in 2019 for LLaMa2, which has a cutoff of September 2022; Touvron et al., 2023b).
@@ 2 TAQA: A QA Dataset for Studying LM Temporal Alignment
To study the temporal chaos of pretrained LMs and how to align them temporally, we curate the TAQA dataset, featuring questions with frequently changing answers in a recent time period (2000-2023). We introduce the formulation of TAQA in §2.1, detail the data construction process in §2.2, and further describe the evaluation metrics in §2.3.
@@ 2.2 Data Construction
To get question $q$ with its multiple temporal answers $(a,t_{s},t_{e})$, we use tables with temporal columns from the Wikipedia dump to build our TAQA dataset, as the temporal columns can be seen as qualifiers $t_{s},t_{e}$ and other columns can be leveraged to generate QA pairs. The advantage of Wikipedia is that its topics are popular, well-structured, and less biased regarding domains. Furthermore, Wikipedia is used in the pretraining corpus for most modern LMs, ensuring that LMs have learned such knowledge from pretraining. We demonstrate our steps for constructing the TAQA dataset as shown in Figure 2 as follows.
@@
Given the Wikipedia dump [^2] We use the English Wikipedia dump of January 1, 2024., we use the WikiExtractor (Attardi, 2015) to extract all tables from the dump into the CSV format. Then, we use heuristic-based methods to identify the columns denoting temporal information (i.e., $(t_{s},t_{e})$) for answers. As we want to gather questions for contemporary knowledge only, we only keep tables with information for every year from 2010 to 2023, resulting in 17,932 tables.
Question generation with LLMs.
@@ 4 Experiments
Model $\downarrow$
Training Cutoff $\downarrow$
$\mathcal{F}^{2022}$ $\mathcal{F}^{2022}_{\text{decay}}$ $\mathcal{F}^{2021}$ $\mathcal{F}^{2021}_{\text{decay}}$ $\mathcal{F}_{\text{max}}$
LLaMA1-65B
Aug. 2022
12.1 26.8 13.2 29.5 54.5
GPT-3
Sep. 2021
10.0 22.9 11.4 25.1 46.5
ChatGPT
Sep. 2021
17.1 32.5 23.8 35.5 46.0
Tülu2-70B
16.2 32.2 19.1 35.1 50.1
LLaMA2-70B
Sep. 2022
17.2 34.1 19.8 37.0 55.9
@@ 6 Related Work
Temporal QA and reasoning are tasks where the correct answers change through time, e.g., “Who is the president of the United States?” Prior research tried using knowledge graphs to tackle these tasks (Jia et al., 2021; Saxena et al., 2021; Shang et al., 2022). As modern LMs can memorize factual knowledge, new benchmarks (Chen et al., 2021; Dhingra et al., 2022; Tan et al., 2023; Wei et al., 2023) address the challenge of leveraging LMs to answer questions with contextual temporal mentions provided (e.g., “in 2023”, “after” certain events, etc.). However, such question structure can sound unnatural, and these studies do not explicitly focus on questions whose answers have changed in recent years. Other news article-based datasets like RealtimeQA (Kasai et al., 2022) and StreamingQA (Liska et al., 2022), despite having more natural questions, still lack questions whose answers can change multiple times, rather than a single time. In contrast, our TAQA dataset curates natural questions without temporal information but with at least five different answers after 2000. Several approaches have been proposed to tackle the temporal QA and reasoning challenges (Son and Oh, 2023; Tan et al., 2023), yet these approaches require dedicated training for enhancing LMs’ understanding of temporal mentions, which are missing in most natural questions.
@@ 7 Conclusion
We build TAQA, a QA dataset with time-sensitive questions during a recent time period (2020-2023). Using this dataset, we quantitatively show that LMs can respond to questions quite chaotically in terms of the time it should be based on. We thus propose the concept of temporal alignment and a set of methods for it. Extensive experiments demonstrate the possibility and effectiveness of aligning model internal knowledge temporally, both to a recent time and to a historical time.
@@ Acknowledgements
We thank labmates at UWNLP for their constructive feedback and intellectual support. This research is supported by Cloud TPUs from Google’s TPU Research Cloud (TRC). We also thank the Beaker team at AI2, who provided the essential computational infrastructure for our experiments.

### v2

@@ Abstract
Language models (LMs) are trained on web text originating from many points in time and, in general, without any explicit temporal grounding. This work investigates the temporal chaos of pretrained LMs and explores various methods to align their internal knowledge to a target time, which we call “temporal alignment.” To do this, we first automatically construct a dataset containing 20K time-sensitive questions and their answers for each year from 2000 to 2023. Based on this dataset, we empirically show that pretrained LMs (e.g., LLaMa2), despite having a recent pretraining cutoff (e.g., 2022), mostly answer questions using earlier knowledge (e.g., in 2019). We then develop several methods, from prompting to finetuning, to align LMs to use their most recent knowledge when answering questions, and investigate various factors in this alignment. Our experiments demonstrate that aligning LLaMa2 to the year 2022 can enhance its performance by up to 62% according to that year’s answers. This improvement occurs even without explicitly mentioning time information, indicating the possibility of aligning models’ internal sense of time after pretraining. Finally, we find that alignment to a historical time is also possible, with up to 2.8×2.8$\times$2.8 × the performance of the unaligned LM in 2010 if finetuning models to that year. These findings hint at the sophistication of LMs'  internal knowledge organization and the necessity of tuning them properly. [^1] Our dataset and code is available at https://github.com/yizhongw/llm-temporal-alignment.
@@
Bowen Zhaothanks: Equal contribution.     Zander Brumbaughfootnotemark:     Yizhong Wangfootnotemark: ♠♣ Hannaneh Hajishirzi♠♣      Noah A. Smith♠♣ University of Washington   Allen Institute for AI bowen98@uw.edu   {brumbzan,yizhongw}@cs.washington.edu
@@ 1 Introduction
We first introduce our TemporalAlignmentQA dataset (TAQA, §2) consisting of 20,148 questions, each with at least five different answers between 2000 and 2023. This dataset is automatically constructed based on Wikipedia tables that contain temporal information, making it easily scalable for future study as the world keeps changing. Complementary to previous temporal QA datasets (Chen et al., 2021; Dhingra et al., 2022a; Kasai et al., 2022; Liska et al., 2022; Tan et al., 2023; Wei et al., 2023), TAQA focuses on facts that changed multiple times during a recent period when most pretraining data is collected, so that it can be used to probe the temporal knowledge distribution of pretrained LMs. Figure 1 plots the F1 score of several representative LMs relative to ground-truth answers from each year, demonstrating that they tend to use earlier knowledge to answer questions, even when they have a very recent pretraining cutoff date (e.g., performance peaks in 2019 for LLaMa2, which has a cutoff of September 2022; Touvron et al., 2023b).
@@ 2 TAQA: A QA Dataset for Studying LM Temporal Alignment
To study the temporal chaos of pretrained LMs and how to align them temporally, we curate the TAQA dataset, featuring questions with frequently changing answers in a recent time period (2000–2023). We introduce the formulation of TAQA in §2.1, detail the data construction process in §2.2, and further describe the evaluation metrics in §2.3.
@@ 2.2 Data Construction
To get question $q$ with its multiple temporal answers $(a,t_{s},t_{e})$, we use tables with temporal columns from the Wikipedia dump to build our TAQA dataset, as the temporal columns can be seen as qualifiers $t_{s},t_{e}$ and other columns can be leveraged to generate QA pairs. The advantage of Wikipedia is that its topics are popular and well-structured. Also, the topic distribution of Wikipedia is less biased than open knowledge graphs, such as Wikidata (Piscopo et al., 2017). Furthermore, Wikipedia is used in the pretraining corpus for most modern LMs, ensuring that LMs have learned such knowledge from pretraining. We demonstrate our steps for constructing the TAQA dataset as shown in Figure 2 as follows.
@@
Given the Wikipedia dump, [^2] We use the English Wikipedia dump of January 1, 2024. we use the WikiExtractor (Attardi, 2015) to extract all tables from the dump into the CSV format. Then, we use heuristic-based methods to identify the columns denoting temporal information (i.e., $(t_{s},t_{e})$) for answers. As we want to gather questions for contemporary knowledge only, we only keep tables with information for every year from 2010 to 2023, resulting in 17,932 tables.
Question generation with LMs.
@@
Train-test split.
Quality control.
To examine the overall quality of the generated TAQA dataset, we randomly sampled 200 questions along with their answers and manually verified their correctness. Among the 200 examples, 185 (92.5%) instances are considered correct, while the incorrectly generated questions are mostly because tables without temporal information are used for QA generation (6/200, 3.0%). Another major issue is that GPT-4 sometimes automatically adds conditional clauses to the question, thus making them not time-sensitive anymore (5/200, 2.5%).
Dataset # Questions
# Post-2019
answer change
Time Median year
Track
length
Avg.
sensitivity
SituatedQA (Zhang and Choi, 2021) 5,061 1,523 1270–2021 2005 8.5 1.7
Templama (Dhingra et al., 2022b) 5,798 1,082 2010–2020 2015 8.9 3.3
TSQA (Chen et al., 2021) 5,061 134 1–2021 1963 26.0 4.4
RealtimeQA (Kasai et al., 2022) 5,152 0 2020–2024 2022 1.0 1.0
TAQA (ours) 20,148 19,316 2000–2023 2013 21.5 15.6
Table 2: Detailed dataset characteristics comparison between TAQA and existing time-sensitive QA datasets. Track length is defined as the average number of years a question’s answer is recorded.
Figure 3: The comparison of TAQA and existing time-sensitive QA datasets regarding the frequency of the answers’ corresponding time. TAQA contains more data in the post-2000 era.
Dataset characteristics
In Table 2, we show a detailed comparison between TAQA and existing time-sensitive QA datasets. Overall, TAQA maintains more time-sensitive questions than existing datasets, making it sufficient for training LMs. Furthermore, questions in TAQA are more sensitive to time than existing datasets, and most answers to questions (95.9%) were changed after 2019 when modern LMs were trained. TAQA also tracks time-sensitive questions over a long period of time (21.5 years on average). Figure 3 also shows that TAQA contains more questions that have corresponding answers recorded after 2000333The Wikipedia tables we use and our intermediate data of TAQA contain answers before 2000 and after 2023, but we discard them for simplicity as these answers are not used in LM temporal alignment analysis.. These properties make TAQA suitable for diagnosing modern LMs’ understanding of time.
@@ 4 Experiments
Model $\downarrow$ Training Cutoff $\downarrow$ $\mathcal{F}^{2022}$ $\mathcal{F}^{2022}_{\text{decay}}$ $\mathcal{F}^{2021}$ $\mathcal{F}^{2021}_{\text{decay}}$ $\mathcal{F}_{\text{max}}$
LLaMA1-65B Aug. 2022 12.1 26.8 13.2 29.5 54.5
GPT-3 Sep. 2021 10.0 22.9 11.4 25.1 46.5
ChatGPT Sep. 2021 17.1 32.5 23.8 35.5 46.0
Tülu2-70B - 16.2 32.2 19.1 35.1 50.1
LLaMA2-70B Sep. 2022 17.2 34.1 19.8 37.0 55.9
@@ 6 Related Work
Temporal QA and reasoning are tasks where the correct answers change through time (Figueroa, 2010), e.g., “Who is the president of the United States”. Prior research tried using knowledge graphs to tackle these tasks (Jia et al., 2021; Saxena et al., 2021; Shang et al., 2022). As modern LMs can memorize factual knowledge, new benchmarks (Chen et al., 2021; Dhingra et al., 2022a; Tan et al., 2023; Wei et al., 2023) address the challenge of leveraging LMs to answer questions with contextual temporal mentions provided (e.g., “in 2023”, “after” certain events, etc.). However, such question structure can sound unnatural, and these studies do not explicitly focus on questions whose answers have changed in recent years. Other news article-based datasets like RealtimeQA (Kasai et al., 2022) and StreamingQA (Liska et al., 2022), despite having more natural questions, still lack questions whose answers can change multiple times, rather than a single time. In contrast, our TAQA dataset curates natural questions without temporal information but with at least five different answers after 2000. Several approaches have been proposed to tackle the temporal QA and reasoning challenges (Son and Oh, 2023; Tan et al., 2023), yet these approaches require dedicated training for enhancing LMs’ understanding of temporal mentions, which are missing in most natural questions.
@@ 7 Conclusion
We build TAQA, a QA dataset with time-sensitive questions during a recent time period (2020–2023). Using this dataset, we quantitatively show that LMs can respond to questions chaotically in terms of the question’s placement in time. We thus propose the concept of temporal alignment and a set of methods for it. Extensive experiments demonstrate the possibility and effectiveness of aligning model internal knowledge temporally, both to a recent time and to a historical time.
@@ Acknowledgements
We thank labmates at UWNLP for their constructive feedback and intellectual support. This work was partially supported by the Office of Naval Research under MURI grant N00014-18-1-2670, and NSF IIS-2044660, an Allen Investigator Distinguished award. This research is also supported by Cloud TPUs from Google’s TPU Research Cloud (TRC) and usage credits from OpenAI. We also thank the Beaker team at AI2, who provided the essential computational infrastructure for our experiments.